Please help me with this one. I had a difficult time trying to do this one. I don't think that I have the right set up. Thanks in advance.


The level of thorium in a sample decreases by a factor of one-half every 2 million years. A meteorite is discovered to have only 8.6% of its original thorium remaining. How old is the meteorite?

Answer by scott8148(6628)

the half-life equation is ___ .086 = (1/2)^(t / 2000000) ___ t is the time (age)

taking logarithm ___ log(.086) = log(1/2) * (t / 2000000)

[log(.086) / log(1/2)] * 2000000 = t = 7,080,000 (approx)